# Cluster algebras and quantum cohomology rings: A-type

### v1

@@ 2.6 Varieties of QPs in $\Omega^{A}_{n}$
    $\sigma_{v_{1}}+\sum_{i=1}^{h}\sigma_{w_{i}}>0,\,\,\sigma_{v_{2}}+\sum_{i=1}^{m}\sigma_{u_{i}}<0\,.$ (2.5)
  2. (2)
    We move to a smaller node $v_{i}$ which is regarded as the frozen node of the subquiver $\mathbf{Q}_{i}$ in Lemma 2.15. Let $\tilde{v}_{1}$ and $\tilde{v}_{2}$ be the two nodes that are smaller than $v_{i}$, and then the conditions for $\sigma_{\tilde{v_{1}}}$ and $\sigma_{\tilde{v}_{2}}$ are similar as (1).
@@
  1. (1)
    If there is only one maximal node that is smaller than $v$, $v_{1}\xrightarrow[]{A_{v_{1}}}v$ or $v_{2}\xleftarrow[]{A_{v_{2}}}v$. Then in the phase described above, $A_{v_{i}}$ $i=1$ or $i=2$ is non-degenerate.
  2. (2)
@@
$Z^{ss}_{\theta}(G)=\{B=0,A_{1}A_{2}=0\,|\,\text{matrices }A_{1},A_{2},A_{3}\text{ nondegenerate}\}.$ (2.6)
@@ 3.1 Gromov-Witten theory and quantum cohomology ring
We refer to the nice book [CK99] about the fundamental properties of Gromov-Witten(GW) theory.
@@
The degree of a stable map $(C,p_{1},\ldots,p_{n};f)$ is defined as the homology class of the image $\beta=f_{*}[C]$. For a fixed curve class $\beta\in H_{2}(\mathcal{X},\mathbb{Z})$, let $\overline{M}_{g,n}(\mathcal{X},\beta)$ denote the stack of stable maps from $n$-marked and genus-g curves $C$ to $\mathcal{X}$ such that $f_{*}[C]=\beta$. When $\mathcal{X}$ is projective, $\overline{M}_{g,n}(\mathcal{X},\beta)$ is a proper separated DM stack and admits a perfect obstruction theory. Hence we can construct the virtual fundamental class $[\overline{M}_{g,n}(\mathcal{X},\beta)]^{vir}\in A_{\text{vdim}}(\overline{M}_{g,n}(\mathcal{X},\beta))$ where $\text{vdim}=\int_{\beta}c_{1}(\mathcal{X})+(\dim(\mathcal{X})-3)(1-g)+n$. See [LT98b, BF97, Beh97].
Let $\pi:\mathcal{C}_{g,n}\rightarrow\overline{M}_{g,n}(\mathcal{X},\beta)\,$ be the universal curve and $s_{i}$ are sections of $\pi$ for each marking $p_{i}$. Let $\omega_{\pi}$ be the relative dualizing sheaf and $\mathcal{P}_{i}=s_{i}^{*}(\omega_{\pi})$ be the cotangent bundle at the $i$-th marking. Define the $\psi$-class by $\psi_{i}:=c_{1}(\mathcal{P}_{i})\in H^{2}(\overline{M}_{g,n}(\mathcal{X},\beta))$. Define evaluation maps by
@@
$\langle\tau_{a_{1}}\gamma_{1},\ldots,\tau_{a_{n}}\gamma_{n}\rangle_{g,n,\beta}^{\mathcal{X}}:=\int_{[\overline{M}_{g,n}(\mathcal{X},\beta)]^{vir}}\prod_{i=1}^{n}\psi_{i}^{a_{i}}ev_{i}^{*}(\gamma_{i})\,.$ (3.2)
@@
$(\gamma_{1}*\gamma_{2},\gamma_{3})=\sum_{\beta}\langle\gamma_{1},\gamma_{2},\gamma_{3}\rangle_{0,3,\beta}^{\mathcal{X}}{q^{\beta}}$ (3.3)
where (,)(,)( , ) is the Poincaré paring of $H^{*}(\mathcal{X})$. Let $R=\mathbb{Q}[[q^{\beta}]]$ be the ring of formal power series in $q^{\beta}$ with coefficients in $\mathbb{Q}$. The ring $H^{*}(\mathcal{X},\mathbb{Q})\otimes R$, equipped with the product *, forms a commutative, associative ring called the small quantum cohomology ring, denoted by $QH^{*}(\mathcal{X})$.
@@
$\langle\tau_{a_{1}}\omega_{1},\ldots,\tau_{a_{n}}\omega_{n}\rangle_{g,n,\beta}^{\mathcal{X},S}=\int_{[\overline{M}_{g,n}(\mathcal{X},\beta)]^{vir,S}}\prod_{i=1}^{n}(\psi_{i}^{S})^{a_{i}}ev_{i}^{*}(\omega_{i})\,$ (3.4)
where $\psi_{i}^{S}\in H^{*}_{S}(\overline{M}_{g,n}(\mathcal{X},\beta))$ are lifts of $\psi_{i}$. Let $F\subset\overline{M}_{g,n}(\mathcal{X},\beta)$ denote a torus fixed loci. There is an induced equivariant perfect obstruction theory on $F$ and
$\langle\tau_{a_{1}}\omega_{1},\ldots,\tau_{a_{n}}\omega_{n}\rangle_{g,n,\beta}^{\mathcal{X},S}=\sum_{F}\int_{F}\frac{i_{F}^{*}\left(\prod_{i=1}^{n}(\psi_{i}^{S})^{a_{i}}ev_{i}^{*}(\omega_{i})\right)}{e^{S}(N^{vir}_{F})}\,.$ (3.5)
The summation is over all torus fixed loci $F$, the map $i_{F}:F\rightarrow\overline{M}_{g,n}(\mathcal{X},\beta)$ is an embedding, and $N_{F}^{vir}$ is the virtual normal bundle of $F$. Suppose $\mathcal{X}$ is projective and $\omega^{\prime}s$ are lifting of $\gamma_{i}^{\prime}s$ in $H^{*}_{S}(\mathcal{X},\mathbb{Q})$. Then the nonequivariant limit of $\langle\tau_{a_{1}}\omega_{1},\ldots,\tau_{a_{n}}\omega_{n}\rangle_{g,n,\beta}^{\mathcal{X},S}$ is equal to the regular Gromov-Witten invariant $\langle\tau_{a_{1}}\gamma_{1},\ldots,\tau_{a_{n}}\gamma_{n}\rangle_{g,n,\beta}^{\mathcal{X}}$. See [GP99].
We can similarly define equivariant quantum product. For $\omega_{1},\omega_{2},\omega_{3}\in H^{*}_{S}(\mathcal{X})$, we define the small quantum product * as follows,
$(\omega_{1}*\omega_{2},\omega_{3})=\sum_{\beta}\langle\omega_{1},\omega_{2},\omega_{3}\rangle^{\mathcal{X},S}_{0,3,\beta}q^{\beta}.$ (3.6)
@@ 3.2 Seiberg duality conjecture
$\theta_{u}^{\prime}=\begin{cases}-\theta_{v}&\text{ for }u=v\\ \theta_{u}+[sign(\sigma_{v})b_{vu}]_{+}\theta_{v}&\text{ for }u\neq v\end{cases}$ (3.7)
@@
$\phi:H^{*}_{S}(\mathcal{Z})\rightarrow H^{*}_{S}(\mathcal{Z}^{\prime}).$ (3.8)
@@
$\int_{\beta}\alpha=\int_{\phi_{*}(\beta)}\phi(\alpha).$ (3.9)
Conjecture 3.3 ([Rua17]).
@@
$<\tau_{a_{1}}\omega_{1},\ldots,\tau_{a_{m}}\omega_{m}>^{\mathcal{Z},S}_{g,n,\beta}=<\tau_{a_{1}}\omega^{\prime}_{1},\ldots,\tau_{a_{m}}\omega_{m}^{\prime}>^{\mathcal{Z}^{\prime},S}_{g,n,\beta^{\prime}}\,.$ (3.10)
@@
$S_{v}^{*}=\begin{cases}S_{v}&\text{if }\sigma_{v}>0\\ S_{v}^{\vee}&\text{ if }\sigma_{v}<0\,.\\ \end{cases}$ (3.11)
@@
$c_{t}^{S}(E)=\sum_{i=0}^{r}(-1)^{i}c_{i}^{S}(E)t^{r-i}$ (3.12)
where $t$ is a formal variable and $c_{i}^{S}(E)$ are the equivariant $i$-th Chern classes. Consider a $\mathbb{Q}$-algebra $QH^{*}_{S}(\mathcal{Z})[t]:=QH^{*}_{S}(\mathcal{Z})\otimes\mathbb{Q}[t]$ which we still call quantum cohomology ring by abuse of terminology. Then Chern polynomials of quotient bundles and tautological bundles $c_{t}^{S}(S_{v})$ and $c_{t}^{S}(Q_{v})$ lie in $QH^{*}_{S}(\mathcal{Z})[t]$. In genus zero, the Seiberg Duality conjecture above implies the following isomorphism which can be viewed as the Seiberg Duality conjecture in quantum cohomology level.
@@
$\Phi:QH^{*}_{S}(\mathcal{Z})[t]\rightarrow QH^{*}_{S}(\mathcal{Z}^{\prime})[t]$ (3.13)
@@ 3.3 Cluster algebra conjecture
$c_{t}^{S}(S_{v})*c_{t}^{S}(Q_{v})=\prod_{v\rightarrow u}c_{t}^{S}(S_{u}^{*})+(-1)^{r_{v}+N_{a}(v)}q_{v}\prod_{w\rightarrow v\in Q_{1}}c_{t}^{S}(S_{w}^{*});$ (3.14)
@@
$c_{t}^{S}(S_{v})*c_{t}^{S}(Q_{v})=\prod_{u\rightarrow v}c_{t}^{S}((S_{u}^{*})^{\vee})+(-1)^{r_{v}+N_{f}(v)}q_{v}^{-1}\prod_{v\rightarrow w\in Q_{1}}c_{t}^{S}((S_{w}^{*})^{\vee}).$ (3.15)
@@
Given a cluster quiver $\mathbf{Q}=(Q_{0},Q_{1},W)$, one can construct a cluster algebra which is called geometric cluster algebra as Section 2.2. Denote the corresponding cluster algebra by $\mathscr{A}_{\mathbf{Q}}$. Let $\mathbf{x}=(x_{u})_{u\in Q_{0}}$ be initial cluster variables.
We introduce another set of formal variables $\{\xi_{i}\}_{i\in Q_{0}}$ and let $q_{v}=(-1)^{N_{f}(v)-N_{a}(v)}\prod_{u\in Q_{0}}\xi_{u}^{b_{uv}}$. Assume that the Conjecture 3.7 holds. Combining the Seiberg duality conjecture and the relation of tautological bundles and quotient bundles in Conjecture 3.6, we propose the following construction for a $\mathbb{Q}$-algebra homomorphism from the cluster algebra to the equivariant Quantum cohomology ring of the associated variety of the decorated QP $QH^{*}_{S}(\mathcal{Z})[t]$.
@@
$\psi:\mathscr{A}_{\mathbf{Q}}\rightarrow QH^{*}_{S}(\mathcal{Z})[t]$ (3.16)
@@
  • $\psi(x_{v}^{\prime})=\begin{cases}(-1)^{N_{f}(v)-r_{v}}\xi_{v}^{-1}\prod_{v\rightarrow u}\xi_{u}c_{t}^{S}(Q_{v})\text{ when }N_{f}(v)>N_{a}(v)\\ (-1)^{N_{a}(v)-r_{v}}\xi_{v}^{-1}\prod_{w\rightarrow v}\xi_{w}c_{t}^{S}(Q_{v})\text{ when }N_{f}(v)<N_{a}(v)\end{cases}$ (3.17)
  • For other non-initial cluster variables $\tilde{x}_{v}$, assume that $\tilde{x}_{v}$ lies in a seed $(\tilde{\mathbf{Q}},\tilde{\mathbf{x}})$ which can be obtained by a sequence of quiver mutations from the initial seed $(\mathbf{Q},\mathbf{x})\xhookrightarrow{\mu_{v_{i}}\cdots\mu_{v_{1}}}(\tilde{\mathbf{Q}},\tilde{\mathbf{x}})$. The conjecture 3.6 asserts that the quantum cohomology rings of varieties of these QPs are isomorphic. Let $\tilde{\Phi}:QH^{*}_{S}(\mathcal{Z})\rightarrow QH^{*}_{S}(\tilde{\mathcal{Z}})$ be the composite of those isomorphisms. Let $\{\tilde{S}_{u}\}_{u\in Q_{0}}$ be tautological bundles over the variety $\tilde{\mathcal{Z}}$ of the decorated QP $(\tilde{\mathbf{Q}},\tilde{\mathbf{r}})$. Then the image of $\tilde{x}_{v}$ is the pullback of $c_{t}^{S}(\tilde{S}_{v})$ if $\tilde{\sigma}_{v}>0$ and pullback of $c_{t}^{S}(\tilde{S}_{v}^{\vee})$ if $\tilde{\sigma}_{v}<0$ scaled by $\xi_{v}$, and those scalars can be determined recursively in order to make $\psi$ a ring homomorphism.
The image of $\tilde{x}_{v}$ for non-initial cluster variables in the last item doesn’t depend on the seed we choose and doesn’t depend on the sequence of quiver mutations we obtain.
One can check that when the Conjecture 3.7 holds, the map $\psi$ defined above preserves all cluster exchange relations.

### v2

@@ 2.6 Varieties of QPs mutation equivalent to A-type
    (5) $\sigma_{v_{1}}+\sum_{i=1}^{h}\sigma_{w_{i}}>0,\,\,\sigma_{v_{2}}+\sum_{i=1}^{m}\sigma_{u_{i}}<0\,.$
  2. 2.
    We move to a smaller node $v_{i}$ which is regarded as the root of the subquiver $\mathbf{Q}_{i}$ in Lemma 2.15. Let $\tilde{v}_{1}$ and $\tilde{v}_{2}$ be the two nodes that are smaller than $v_{i}$, and then the conditions for $\sigma_{\tilde{v_{1}}}$ and $\sigma_{\tilde{v}_{2}}$ are similar as (1).
@@
  1. 1.
    If there is only one maximal node that is smaller than $v$, $v_{1}\xrightarrow[]{A_{v_{1}}}v$ or $v_{2}\xleftarrow[]{A_{v_{2}}}v$, then in the phase described above, $A_{v_{i}}$ $i=1$ or $i=2$ is non-degenerate.
  2. 2.
@@
$Z^{ss}_{\theta}(G)=\{B=0,A_{1}A_{2}=0\,|\,\text{matrices }A_{1},A_{2},A_{3}\text{ nondegenerate}\}.$
@@ 3.1 Gromov-Witten theory and quantum cohomology ring
We refer to the nice book [GW:mirrorsym] about the fundamental properties of Gromov-Witten(GW) theory.
@@
The degree of a stable map $(C,p_{1},\ldots,p_{n};f)$ is defined as the homology class of the image $\beta=f_{*}[C]$. For a fixed curve class $\beta\in H_{2}(\mathcal{X},\mathbb{Z})$, let $\overline{M}_{g,n}(\mathcal{X},\beta)$ denote the stack of stable maps from the genus-g curves $C$ with $n$ markings to $\mathcal{X}$ such that $f_{*}[C]=\beta$. When $\mathcal{X}$ is projective, $\overline{M}_{g,n}(\mathcal{X},\beta)$ is a proper and separated DM stack and admits a perfect obstruction theory. Hence, we can construct the virtual fundamental class $[\overline{M}_{g,n}(\mathcal{X},\beta)]^{vir}\in A_{\text{vdim}}(\overline{M}_{g,n}(\mathcal{X},\beta))$ where $\text{vdim}=\int_{\beta}c_{1}(\mathcal{X})+(\dim(\mathcal{X})-3)(1-g)+n$. See [GW:LT, GW:BFintrinsic, GW:B].
Let $\pi:\mathcal{C}_{g,n}\rightarrow\overline{M}_{g,n}(\mathcal{X},\beta)\,$ be the universal curve and $s_{i}$ be sections of $\pi$ for each marking $p_{i}$. Let $\omega_{\pi}$ be the relative dualizing sheaf and $\mathcal{P}_{i}=s_{i}^{*}(\omega_{\pi})$ be the cotangent bundle at the $i$-th marking. Define the $\psi$-class by $\psi_{i}:=c_{1}(\mathcal{P}_{i})\in H^{2}(\overline{M}_{g,n}(\mathcal{X},\beta))$. Define evaluation maps by
@@
(6) $\langle\tau_{a_{1}}\gamma_{1},\ldots,\tau_{a_{n}}\gamma_{n}\rangle_{g,n,\beta}^{\mathcal{X}}:=\int_{[\overline{M}_{g,n}(\mathcal{X},\beta)]^{vir}}\prod_{i=1}^{n}\psi_{i}^{a_{i}}ev_{i}^{*}(\gamma_{i})\,.$
@@
$(\gamma_{1}*\gamma_{2},\gamma_{3})=\sum_{\beta}\langle\gamma_{1},\gamma_{2},\gamma_{3}\rangle_{0,3,\beta}^{\mathcal{X}}{q^{\beta}}$
where (,)(,)( , ) is the Poincaré pairing of $H^{*}(\mathcal{X})$. Let $R=\mathbb{Q}[[q^{\beta}]]$ be the ring of formal power series in $q^{\beta}$ with coefficients in $\mathbb{Q}$. The ring $H^{*}(\mathcal{X},\mathbb{Q})\otimes R$, equipped with the product *, forms a commutative associative ring called the small quantum cohomology ring, denoted by $QH^{*}(\mathcal{X})$.
@@
$\langle\tau_{a_{1}}\omega_{1},\ldots,\tau_{a_{n}}\omega_{n}\rangle_{g,n,\beta}^{\mathcal{X},S}=\int_{[\overline{M}_{g,n}(\mathcal{X},\beta)]^{vir,S}}\prod_{i=1}^{n}(\psi_{i}^{S})^{a_{i}}ev_{i}^{*}(\omega_{i})\,$
where $\psi_{i}^{S}\in H^{*}_{S}(\overline{M}_{g,n}(\mathcal{X},\beta))$ are lifts of $\psi_{i}$. Let $F\subset\overline{M}_{g,n}(\mathcal{X},\beta)$ denote a torus-fixed locus. There is an induced equivariant perfect obstruction theory on $F$ and
$\langle\tau_{a_{1}}\omega_{1},\ldots,\tau_{a_{n}}\omega_{n}\rangle_{g,n,\beta}^{\mathcal{X},S}=\sum_{F}\int_{F}\frac{i_{F}^{*}\left(\prod_{i=1}^{n}(\psi_{i}^{S})^{a_{i}}ev_{i}^{*}(\omega_{i})\right)}{e^{S}(N^{vir}_{F})}\,.$
The summation is over all torus fixed loci $F$, the map $i_{F}:F\rightarrow\overline{M}_{g,n}(\mathcal{X},\beta)$ is an embedding, and $N_{F}^{vir}$ is the virtual normal bundle of $F$. Suppose $\mathcal{X}$ is projective and $\omega^{\prime}s$ are lifting of $\gamma_{i}^{\prime}s$ in $H^{*}_{S}(\mathcal{X},\mathbb{Q})$. Then the nonequivariant limit of $\langle\tau_{a_{1}}\omega_{1},\ldots,\tau_{a_{n}}\omega_{n}\rangle_{g,n,\beta}^{\mathcal{X},S}$ is equal to the regular Gromov-Witten invariant $\langle\tau_{a_{1}}\gamma_{1},\ldots,\tau_{a_{n}}\gamma_{n}\rangle_{g,n,\beta}^{\mathcal{X}}$. See [GW:GPequiv].
We can similarly define an equivariant quantum product. For $\omega_{1},\omega_{2},\omega_{3}\in H^{*}_{S}(\mathcal{X})$, we define the small quantum product * as follows,
$(\omega_{1}*\omega_{2},\omega_{3})=\sum_{\beta}\langle\omega_{1},\omega_{2},\omega_{3}\rangle^{\mathcal{X},S}_{0,3,\beta}q^{\beta}.$
@@ 3.2 Seiberg duality conjecture
(7) $\theta_{u}^{\prime}=\begin{cases}-\theta_{v}&\text{ for }u=v\\ \theta_{u}+[sign(\sigma_{v})b_{vu}]_{+}\theta_{v}&\text{ for }u\neq v\end{cases}$
@@
$\phi:H^{*}_{S}(\mathcal{Z})\rightarrow H^{*}_{S}(\mathcal{Z}^{\prime}).$
@@
(8) $\int_{\beta}\alpha=\int_{\phi_{*}(\beta)}\phi(\alpha).$
Conjecture 3.3 ([nonabelianGLSM:YR]).
@@
$<\tau_{a_{1}}\omega_{1},\ldots,\tau_{a_{m}}\omega_{m}>^{\mathcal{Z},S}_{g,n,\beta}=<\tau_{a_{1}}\omega^{\prime}_{1},\ldots,\tau_{a_{m}}\omega_{m}^{\prime}>^{\mathcal{Z}^{\prime},S}_{g,n,\beta^{\prime}}\,.$
@@
$S_{v}^{*}=\begin{cases}S_{v}&\text{if }\sigma_{v}>0\\ S_{v}^{\vee}&\text{ if }\sigma_{v}<0\,.\\ \end{cases}$
@@
(9) $c_{t}^{S}(E)=\sum_{i=0}^{r}(-1)^{i}c_{i}^{S}(E)t^{r-i}$
where $t$ is a formal variable and $c_{i}^{S}(E)$ are the equivariant $i$-th Chern classes. Consider a $\mathbb{Q}$-algebra $QH^{*}_{S}(\mathcal{Z})[t]:=QH^{*}_{S}(\mathcal{Z})\otimes\mathbb{Q}[t]$ which we still call quantum cohomology ring by abuse of terminology. Then the Chern polynomials of the quotient bundles and tautological bundles $c_{t}^{S}(S_{v})$ and $c_{t}^{S}(Q_{v})$ lie in $QH^{*}_{S}(\mathcal{Z})[t]$. In genus zero, the Seiberg Duality conjecture above implies the following isomorphism, which can be viewed as the Seiberg Duality conjecture at the quantum cohomology level.
@@
$\Phi:QH^{*}_{S}(\mathcal{Z})[t]\rightarrow QH^{*}_{S}(\mathcal{Z}^{\prime})[t]$
@@ 3.3 Cluster algebra conjecture
(10) $c_{t}^{S}(S_{v})*c_{t}^{S}(Q_{v})=\prod_{v\rightarrow u}c_{t}^{S}(S_{u}^{*})+(-1)^{r_{v}+N_{a}(v)}q_{v}\prod_{w\rightarrow v\in Q_{1}}c_{t}^{S}(S_{w}^{*});$
@@
(11) $c_{t}^{S}(S_{v})*c_{t}^{S}(Q_{v})=\prod_{u\rightarrow v}c_{t}^{S}((S_{u}^{*})^{\vee})+(-1)^{r_{v}+N_{f}(v)}q_{v}^{-1}\prod_{v\rightarrow w\in Q_{1}}c_{t}^{S}((S_{w}^{*})^{\vee}).$
@@
Given a cluster quiver $\mathbf{Q}=(Q_{0},Q_{1},W)$, one can construct a cluster algebra which is called a geometric cluster algebra as in Section 2.2. Denote the corresponding cluster algebra by $\mathscr{A}_{\mathbf{Q}}$. Let $\mathbf{x}=(x_{u})_{u\in Q_{0}}$ be the initial cluster variables.
We introduce another set of formal variables $\{\xi_{i}\}_{i\in Q_{0}}$ and let $q_{v}=(-1)^{N_{f}(v)-N_{a}(v)}\prod_{u\in Q_{0}}\xi_{u}^{b_{uv}}$. Assume that Conjecture 3.7 holds. Combining the Seiberg duality conjecture and the relation of tautological bundles and quotient bundles in Conjecture 3.6, we propose the following construction for a $\mathbb{Q}$ algebra homomorphism from the cluster algebra to the equivariant quantum cohomology ring of the associated variety of decorated QP $QH^{*}_{S}(\mathcal{Z})[t]$.
@@
$\psi:\mathscr{A}_{\mathbf{Q}}\rightarrow QH^{*}_{S}(\mathcal{Z})[t]$
@@
  2. 2.
    $\psi(x_{v}^{\prime})=\begin{cases}(-1)^{N_{f}(v)-r_{v}}\xi_{v}^{-1}\prod_{v\rightarrow u}\xi_{u}c_{t}^{S}(Q_{v})\text{ when }N_{f}(v)>N_{a}(v)\\ (-1)^{N_{a}(v)-r_{v}}\xi_{v}^{-1}\prod_{w\rightarrow v}\xi_{w}c_{t}^{S}(Q_{v})\text{ when }N_{f}(v)<N_{a}(v)\end{cases}$
  3. 3.
    For other non-initial cluster variables $\tilde{x}_{v}$, assume that $\tilde{x}_{v}$ lies in a seed $(\tilde{\mathbf{Q}},\tilde{\mathbf{x}})$ that can be obtained by a sequence of quiver mutations from the initial seed $(\mathbf{Q},\mathbf{x})\xhookrightarrow{\mu_{v_{i}}\cdots\mu_{v_{1}}}(\tilde{\mathbf{Q}},\tilde{\mathbf{x}})$. The conjecture 3.6 asserts that the quantum cohomology rings of varieties of these QPs are isomorphic. Let $\tilde{\Phi}:QH^{*}_{S}(\mathcal{Z})\rightarrow QH^{*}_{S}(\tilde{\mathcal{Z}})$ be the composite of these isomorphisms. Let $\{\tilde{S}_{u}\}_{u\in Q_{0}}$ be tautological bundles over the variety $\tilde{\mathcal{Z}}$ of the decorated QP $(\tilde{\mathbf{Q}},\tilde{\mathbf{r}})$. Then the image of $\tilde{x}_{v}$ is the pullback of $c_{t}^{S}(\tilde{S}_{v})$ if $\tilde{\sigma}_{v}>0$ and the pullback of $c_{t}^{S}(\tilde{S}_{v}^{\vee})$ if $\tilde{\sigma}_{v}<0$ scaled by $\xi_{v}$, and these scalars can be determined recursively to make $\psi$ a ring homomorphism.
The image of $\tilde{x}_{v}$ for noninitial cluster variables in the last item does not depend on the seed we choose and does not depend on the sequence of quiver mutations we obtain.
One can check that when Conjecture 3.7 holds, the map $\psi$ defined above preserves all cluster exchange relations.